A preventive maintenance program has been conducted to replace a polymer gear with a required functional stress level of 20 MPa every two years. The polymer material was tested to have an initial stress of 36 MPa, which was dropped to 30 MPa after a year usage at a constant strain. Provide your engineering reasoning whether the maintenance program for the polymer gear with respect to reliability and cost is correct or not.
Answer #1

bata from the girco question 60 = 36mpa - 30 mpa t =1 (12 months we know year 4) GoC - 0 where 6 = stress impa) Go - Original(-24166) 20 - Gà C (-0.3636) 20 - 60 C 20 =GO - (-0.3636) e 20 0.695 28.77 :: Go = 28.79 mpa initial stress 93 still mainted